OpenStudy (gahm8684):

Just need someone to check my answer or to explain me if im wrong. Giving medals. thanks(:.

OpenStudy (gahm8684):

The distance traveled, in feet, of a ball dropped from a tall building is modeled by the equation d(t) = 16t2 where d equals the distance traveled at time t seconds and t equals the time in seconds. What does the average rate of change of d(t) from t = 2 to t = 5 represent? The ball travels an average distance of 112 feet from 2 seconds to 5 seconds. The ball falls down with an average speed of 48 feet per second from 2 seconds to 5 seconds. The ball falls down with an average speed of 112 feet per second from 2 seconds to 5 seconds. The ball travels an average distance of 48 feet from 2 seconds to 5 seconds.

OpenStudy (gahm8684):

I say its a

OpenStudy (gahm8684):

am a 84 percent sure it is a if not its c

OpenStudy (gahm8684):

help please guys:) @JoannaBlackwelder

OpenStudy (gahm8684):

@lankymonkey

OpenStudy (gahm8684):

@nikato

OpenStudy (anonymous):

A :) Good job!!!!

OpenStudy (anonymous):

Don't forget to pick Best response. Bye.

OpenStudy (anonymous):

The answer is A. =)

OpenStudy (anonymous):

lol

OpenStudy (gahm8684):

this is my work: d(2)=16(2)^2 d(2)=16(4) d(2)=64 d(5)=16(5)^2 d(5)=16(25) d(5)=400 400-64/5-2=112

OpenStudy (gahm8684):

thanks

OpenStudy (anonymous):

welcome!!! :)

OpenStudy (gahm8684):

i got more ill tag you. thanks guys forcoming:)

OpenStudy (anonymous):

kk

OpenStudy (joannablackwelder):

I don't think it is A. The distance traveled would be 400-64 ft.

OpenStudy (joannablackwelder):

And what does that have to do with the rate of change?

OpenStudy (gahm8684):

is it c then?;o

OpenStudy (joannablackwelder):

Yeah, that is what I think.

OpenStudy (joannablackwelder):

change in distance/change in time :)

OpenStudy (gahm8684):

ooouuup thanks

OpenStudy (joannablackwelder):

No worries

OpenStudy (gahm8684):

youre really smart are you a teacher?

OpenStudy (joannablackwelder):

Thanks. I'm a tutor. :)

OpenStudy (gahm8684):

i knew it

OpenStudy (gahm8684):

youre really good explaining

OpenStudy (joannablackwelder):

Thanks! I try. :)